OpenStudy (anonymous):

The manufacturing cost of a cell phone was $124. Sellers first marked up the cost by 20 percent. Sales were poor, so sellers decreased the markup to 15 percent. Calculate the old and new selling prices.

OpenStudy (anonymous):

20% = 0.2 so the old price would be $124* 0.2= Mark Up So old price would be $124+Mark up Could you solve for it?

OpenStudy (campbell_st):

well the 20% increase means you pay P = 124 * (1+0.2) as the markup price then discounting you pay D = P*(1 - 0.15) hope it helps